# Complements of discriminants of real parabolic function singularities. II

### 08 Mar 2026 v5

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
Components with $\mbox{\rm Card}=1648$. For each polynomial $f$ of class $\Xi_{2}$, the polynomials fCf-C with sufficiently large $|C|$ and different signs of $C$ realize two different virtual components with $\mbox{Ind}=0$. The homology indices of these components are equal to $(000)$ and $(011)$. For example, let $f$ be the $S(3)$-invariant polynomial
@@
has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2. Thus, it has $\mbox{Ind}=1$ and homology index $(001)$. This polynomial is invariant under the involution (45). According to the last statement of Proposition 42, the associated virtual component cannot have $\mbox{Card}=94$. Therefore, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
@@
The group $H_{*}(W_{0},W_{-})$ corresponding to any polynomial fCf-C with $C\in\left(0,\frac{88}{49}\right)$ is isomorphic to $\mathbb{Z}$} in dimension 0 and is trivial in all other dimensions. Therefore, the homology index $\mbox{HI}(f-1)$ is $(100)$. The involution (45) moves this polynomial to one with $\mbox{HI}=(012)$. In particular, these two polynomials represent two different components. They are $S(3)$-invariant. Therefore, by Proposition 35, the associated virtual components are also different. According to Proposition 42 and Table 3, these are two components with $\mbox{Card}=94$.
According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(020)$ or $(130)$, while its image under the involution (45) has the homology index $(031)$. Therefore, these two polynomials belong to different components with $\mbox{Ind}=-2$ and equal Card invariants. The only possibility for this is two components with $\mbox{Card}=336.$
@@
Figure 7. Coxeter–Dynkin graph for (58)
@@
Figure 8. Coxeter-Dynkin graph for (59)
@@
According to Corollary 34, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data. According to Proposition 41 of [29], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2. Therefore the polynomial $f_{0}$ has no local minima or maxima.
Let $s:[0,1]\to{\mathbb{R}}^{8}$ be the parametrized line segment such that $s(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $s(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$. Using the Lyashko–Looijenga covering (whose existence for the spaces $\Xi_{2}$ is shown in [12]) one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $s(t)$ is the sequence of critical values of the polynomial $I(t)$. As in the proof of Theorem 10 of [29], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined. Indeed, according to Proposition 2 of [12], this is guaranteed by the fact that all intersection forms corresponding to the limit positions $c_{-}$ and $c_{+}$ of the critical values of the polynomials $I(t)$ are elliptic. (These forms are given respectively by the upper left and the bottom right $3\times 3$ corners of the matrix (59).) The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with critical value $c_{-}$ and three critical points with critical value $c_{+}$.
@@
By statement 3 of Proposition 42, the involution (45) moves the polynomial $I(1)$ to a polynomial in the different component of the discriminant complement with $\mbox{Card}=60$. Thus, this component is also represented by an $S(3)$-invariant polynomial. This proves the assertion of Theorem 43 for both virtual components with $\mbox{Card}=60$.
@@
  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $\varepsilon^{2}$,
@@
Denote the polynomial $f_{\varepsilon}-1$ by $\bar{f}$. It belongs to the same component of the discriminant complement as the polynomial $\hat{f}-1$. The canonical system of paths in ${\mathbb{C}}^{1}$ defining the frame of vanishing cycles in $H_{2}(\bar{f}_{\varepsilon}^{-1}(0))$ then looks as shown in Fig. 9 (left), where the numbers 1–8 denote the order of basic vanishing cycles corresponding to the critical values.
@@
The virtual function of the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program. Based on the information on the polynomial $\bar{f}$, let us try to find it in this list. According to § 5.2 of [29] there are exactly 1897 virtual functions of class $\Xi_{2}$ with six real critical points. Only 140 of these virtual functions are associated with real polynomials that have a minimum point. Each of these haves, in addition to the minimum, three critical points with Morse index 1 and two points with Morse index 2. Exactly $20=140/7$ of these virtual functions have four negative and two positive critical values at real points. Of these 20 virtual functions, only 12 have all these positive critical values at the critical points with Morse index 2. Thus, the virtual function associated with $\bar{f}$ is among these 12. Upon a special request, our program printed these 12 virtual functions. It turns out that for all of them the intersection index $\langle\Delta_{7},\Delta_{8}\rangle$ of two cycles vanishing at non-real points is equal to 1. In particular, this is true for the real polynomial $\bar{f}$. Using the Lyashko–Looijenga map in the form of [12] (see the proof of Theorem 7 in [29]), we can move this polynomial within the discriminant complement in $\Xi_{2}$ in such a way that all its real critical values stay fixed, but two non-real critical values move along the corresponding paths shown in Fig. 9 (left) almost until the point 0, then travel close to the real axis to the point $1\in{\mathbb{R}}$ and finally meet at this point. Due to the equality $\langle\Delta_{7},\Delta_{8}\rangle=1,$ the critical points corresponding to these values also meet at a real critical point of class $A_{2}$. With a slight further perturbation, we obtain a generic polynomial with eight real critical points, whose six critical values are the same as those $\bar{f}$, and the other two critical points, obtained from the $A_{2}$ point, have real critical values $1-\zeta$ and $1+\zeta$ with some very small $\zeta>0$. Denote this polynomial by $f!$, see Fig. 9 (right). Its critical point corresponding to the critical value $\varepsilon-1$ and marked with 1 in this figure is a minimum point. The critical points with values marked with 2, 3, and 4 have Morse index 1, and the critical points marked with 7 and 8 have Morse index 2. According to Lemmas V.3.2 and V.3.3 of [23], the newborn critical point with critical value $1-\zeta$ (marked with 5 in the picture) has an odd Morse index, and the point with critical value $1+\zeta$ has an even Morse index. According to Proposition 43 of [29], a generic polynomial of class $\Xi_{2}$ with eight real critical points cannot have more than one point of local extremum, therefore these odd and even Morse indices are in fact equal to one and two, respectively.
@@
Proposition 43 of [29] implies that all real polynomials of class $\Xi_{2}$ with eight critical points, one of which is a minimum, have the Coxeter-Dynkin graph shown in Fig. 7, in such a way that
@@
The virtual function associated with the polynomial $f!$ is determined by this graph supplied with the information on the Morse indices of critical points corresponding to its vertices, and with the numbering of these vertices corresponding to the ascending order of critical values, i.e., the correspondence of these vertices with the black points in Fig. 9 right. The graph and the Morse indices are known to us, while concerning the numbering we know the following. The lowest vertex in Fig. 7 has number 1, the vertices in the middle row have numbers 2, 3, 4, and 5, and the vertices in the upper row have the numbers 6, 7, and 8. The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5. Indeed, the intersection index $\langle\Delta_{5},\Delta_{6}\rangle$ of cycles vanishing in two newborn Morse critical points is equal to 1, see Lemma V.3.2 in [23]. In particular, the positive critical values of the polynomial $f!$ correspond to all vertices in the top row of Fig. 7 and one vertex of valence three in the middle row.
All virtual functions with intersection matrix defined by the graph of Fig. 7 and the Morse indices and numbering of vertices satisfying these restrictions belong to the same virtual component. Indeed, they can all be obtained from each other by transpositions of orders of vertices that are not connected by edges of the graph and correspond to the critical values of the same sign. Such a transposition is a flip of virtual functions which corresponds to the elementary function surgery when one critical value overtakes another critical value at a distant critical point. Our program, when applied to one of the virtual functions from this set, says that this virtual component has 1216 vertices. $\Box$
@@
The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l). Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin. These numbers $c(l)$ form a lower semi-continuous function on $L/{\mathbb{Z}}_{2}$. By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$. The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant. This union is connected and invariant under the action (44) of the group $S(3)$. By Lemma 48, for $l\equiv x-y$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$. The polynomial $F(l)-\delta$ also belongs to this component because all the polynomials F(l)tF(l)-t, where $t\in[\delta,1]$, are non-discriminant. Therefore, the entire component is $S(3)$-invariant. This concludes the proof of Theorem 43 for the virtual component with $\mbox{Card}=1216$.
Components with $\mbox{Card}=1318$ and $844$. Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other real critical points with critical value 0. By the versality of the deformation (42), this stratum is a smooth submanifold of codimension four in $\Xi_{2}$. All polynomials $F(l)$ of the form (61) belong to this stratum.
Consider a tubular neighborhood of this stratum in $\Xi_{2}$ fibered into its orthogonal slices diffeomorphic to small four-dimensional balls centered at the points of the stratum. Suppose it is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is a versal deformation of the critical point of the function $F(l)$ at the origin. Therefore, all the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
(62) $x^{2}y-y^{3}+\varkappa_{1}+\varkappa_{2}x+\varkappa_{3}y+\varkappa_{4}y^{2}$
of the sample polynomial $x^{2}y-y^{3}$ of class $D_{4}^{-}$ in two variables with parameters $\varkappa_{i}$.
Denote the fiber of this tubular neighborhood over the point $\{F(l)\}$, $l\in L$, by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
@@
    $f$ has four real critical points near the origin obtained by splitting the critical point of the polynomial $F(l)$: specifically, one minimum and three points with Morse index 1,
@@
Similarly, define the subset $W(l)\subset U(l)$ as the union of points satisfying condition (a) but with condition (b) replaced by $\mbox{Card}(f)=1318$.
@@
Preliminary remarks. By the versality property, the discriminant is equisingular along the smooth stratum $\{D_{4}^{-}\}$. Therefore, it is sufficient to prove Lemma 50 for only one particular $l\in L$, say for $l\equiv x-y$.
@@
To prove Lemma 50, we will describe the geometry of the discriminant in the parameter space of the versal deformation (62) in detail. The Morse polynomials of the form (62) can have two or four critical points depending on the parameters $\varkappa_{2},\varkappa_{3},$ and $\varkappa_{4}$. In a neighborhood of the origin of the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse functions looks as shown in Fig. 11 (left), see [3], § 21.3.
@@
This set divides the space ${\mathbb{R}}^{3}$ into three parts. The larger part consists of polynomials with only two real critical points. Two symmetric “pyramids” are filled with polynomials (62) that have three saddlepoints and one critical point that is either a minimum or a maximum, depending on the pyramid. We are only interested in the polynomials (62) with a minimum point. For each generic interior point $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ of the corresponding pyramid, the set of non-discriminant polynomials (62) with these values consists of five intervals. Indeed, the coefficient $\varkappa_{1}$ of a polynomial from this set should be not equal to minus a critical value of the polynomial
(63) $x^{2}y-y^{3}+\varkappa_{2}x+\varkappa_{4}y+\varkappa_{4}y^{2},$
@@
i.e., it can belong to one of the five intervals into which these critical values divide the line $\mathbb{R}$}. For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal. The set of these non-generic polynomials (63) divides the pyramid into six parts. The intersection of the pyramid with a plane perpendicular to its axis looks as the curvilinear triangle shown in Fig. 11 (center). The non-generic polynomials form three segments in it. In particular, the central point of this triangle corresponds to a polynomial $f_{\triangle}$ with equal critical values at all three saddlepoints. The zero-level set of the function $f_{\triangle}-\tau$, where $\tau$ is the critical value, looks as in Fig. 10 (right). These polynomials, representing various sections of the pyramid and corresponding to different values of $\tau$, sweep out its central axis. An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
@@
Let us label the three saddlepoints of an arbitrary polynomial inside the pyramid as $a,b,c$. This labeling of critical points can be uniquely extended by continuity to all other polynomials within the pyramid. Each of the six cells, into which the curvilinear triangle in Fig. 11 (center) is split by the segments, is characterized by the order in $\mathbb{R}$} of the critical values of the corresponding polynomials at the marked points.

### 16 Mar 2026 v6

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
Components with $\mbox{\rm Card}=1648$. For each polynomial $f$ of class $\Xi_{2}$, the polynomials fCf-C with sufficiently large $|C|$ and different signs of $C$ realize two different virtual components with $\mbox{Ind}=0$. The homology indices of these components are equal to $(0,0,0)$ and $(0,1,1)$. For example, let $f$ be the $S(3)$-invariant polynomial
@@
has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2. Thus, it has $\mbox{Ind}=1$ and homology index $(0,0,1)$. This polynomial is invariant under the involution (45). According to the last statement of Proposition 43, the associated virtual component cannot have $\mbox{Card}=94$. Therefore, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
@@
The group $H_{*}(W_{0},W_{-})$ corresponding to any polynomial fCf-C with $C\in\left(0,\frac{88}{49}\right)$ is isomorphic to $\mathbb{Z}$} in dimension 0 and is trivial in all other dimensions. Therefore, the homology index $\mbox{HI}(f-1)$ is $(1,0,0)$. The involution (45) transforms this polynomial into one with $\mbox{HI}=(0,1,2)$. In particular, these two polynomials represent two different components. They are $S(3)$-invariant. The associated virtual components are therefore associated with single real connected components and are also different, though their Card invariants are the same. According to Proposition 43 and Table 3, these are two components with $\mbox{Card}=94$.
According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(0,2,0)$ or $(1,3,0)$, while its image under the involution (45) has the homology index $(0,3,1)$. Therefore, these two polynomials belong to different components with $\mbox{Ind}=-2$ and equal Card invariants. The only possibility for this is two components with $\mbox{Card}=336.$
@@
Figure 7. Coxeter–Dynkin graph for the polynomial (58)
@@
Figure 8. Coxeter-Dynkin graph of the matrix (59)
@@
According to Corollary 35, there exists a real polynomial $f_{0}\in\Xi_{2}$ with these data. According to Proposition 41 of [30], for all functions in class $\Xi_{2}$ that satisfy these conditions 1)–4), the odd Morse indices of their critical points are all equal to 1, and the even Morse indices are all equal to 2. Therefore, the polynomial $f_{0}$ has no local minima or maxima.
Let $s:[0,1]\to{\mathbb{R}}^{8}$ be the parameterized line segment such that $s(0)$ is the collection of critical values $(c_{1},\dots,c_{8})$ of the polynomial $f_{0}$, ordered by increasing value, and $s(1)=(c_{-},c_{-},c_{-},c_{4},c_{5},c_{+},c_{+},c_{+})$, where $c_{-}$ is the mean value of $c_{1},c_{2},$ and $c_{3}$, and $c_{+}$ is the mean value of $c_{6},c_{7},$ and $c_{8}$. Using the Lyashko–Looijenga covering, whose existence in the spaces $\Xi_{2}$ is shown in [12], one can uniquely construct a path $I:[0,1)\to\Xi_{2}$ such that $I(0)=f_{0}$ and, for any $t\in[0,1)$, $s(t)$ is the sequence of critical values of the polynomial $I(t)$. As in the proof of Theorem 10 of [30], the limit point $I(1)\in\Xi_{2}$ of this lifted path is well-defined. Indeed, according to Proposition 2 of [12], this is guaranteed by the fact that all intersection forms corresponding to the limit positions $c_{-}$ and $c_{+}$ of the critical values of the polynomials $I(t)$ are elliptic. (These forms are given respectively by the upper left and the bottom right $3\times 3$ corners of the matrix (59).) The point $I(1)$ is a Morse (albeit non-generic) polynomial of class $\Xi_{2}$ having three critical points with critical value $c_{-}$ and three critical points with critical value $c_{+}$.
@@
By statement 3 of Proposition 43, the involution (45) moves the polynomial $I(1)$ to a polynomial in a different component of the discriminant complement with $\mbox{Card}=60$. Thus, this component is also represented by an $S(3)$-invariant polynomial. This proves the assertion of Theorem 44 for both virtual components with $\mbox{Card}=60$.
@@
  • its critical point of class $D_{4}^{-}$ splits into one minimum point with critical value $\varepsilon$, and three critical points with Morse index 1 and critical values 0, $\varepsilon^{2},$ and $\varepsilon^{2}$;
@@
Denote the polynomial $f_{\varepsilon}-1$ by $\bar{f}$. It belongs to the same component of the discriminant complement as the polynomial $\hat{f}-1$. The canonical system of paths in ${\mathbb{C}}^{1}$ defining the frame of vanishing cycles in $H_{2}(\bar{f}^{-1}(0))$ then looks as shown in Fig. 9 (left), where the numbers 1–8 denote the order of basic vanishing cycles corresponding to the critical values.
@@
The virtual function associated with the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program. Based on the information on the polynomial $\bar{f}$, let us try to find it in this list. According to § 5.2 of [30] there are exactly 1897 virtual functions of class $\Xi_{2}$ with six real critical points. Only 140 of these virtual functions are associated with real polynomials that have a minimum point. Each of these haves, in addition to the minimum, three critical points with Morse index 1 and two points with Morse index 2. Exactly $20=140/7$ of these virtual functions have four negative and two positive critical values at real points. Of these 20 virtual functions, only 12 have all these positive critical values at the critical points with Morse index 2. Thus, the virtual function associated with $\bar{f}$ is among these 12. Upon a special request, our program printed these 12 virtual functions. It turns out that for all of them the intersection index $\langle\Delta_{7},\Delta_{8}\rangle$ of two cycles vanishing at non-real points is equal to 1. In particular, this is true for the real polynomial $\bar{f}$. Using the Lyashko–Looijenga map in the form of [12] (see the proof of Theorem 7 in [30]), we can move this polynomial within the discriminant complement in $\Xi_{2}$ in such a way that all its real critical values stay fixed, but two non-real critical values move along the corresponding paths shown in Fig. 9 (left) almost until the point 0, then travel close to the real axis to the point $1\in{\mathbb{R}}$ and finally meet at this point. Due to the equality $\langle\Delta_{7},\Delta_{8}\rangle=1,$ the critical points corresponding to these values also meet at a real critical point of class $A_{2}$. With a slight further perturbation, we obtain a generic polynomial with eight real critical points, whose six critical values are the same as those $\bar{f}$, and the other two critical points, obtained from the $A_{2}$ point, have real critical values $1-\zeta$ and $1+\zeta$ with some very small $\zeta>0$. Denote this polynomial by $f!$, see Fig. 9 (right). Its critical point corresponding to the critical value $\varepsilon-1$ and marked with 1 in this figure is a minimum point. The critical points with values marked with 2, 3, and 4 have Morse index 1, and the critical points marked with 7 and 8 have Morse index 2. According to Lemmas V.3.2 and V.3.3 of [24], the newborn critical point with critical value $1-\zeta$ (marked with 5 in the picture) has an odd Morse index, and the point with critical value $1+\zeta$ has an even Morse index. According to Proposition 43 of [30], a generic polynomial of class $\Xi_{2}$ with eight real critical points cannot have more than one point of local extremum, therefore these odd and even Morse indices are in fact equal to one and two, respectively.
@@
Proposition 43 of [30] implies that the intersection form of any real polynomial of class $\Xi_{2}$ with eight critical points, one of which is a minimum, is defined by the Coxeter-Dynkin graph shown in Fig. 7, in such a way that
@@
The virtual function associated with the polynomial $f!$ is determined by this graph supplied with the information on the Morse indices of critical points corresponding to its vertices, and with the numbering of these vertices corresponding to the ascending order of critical values, i.e., the correspondence of these vertices with the black points in Fig. 9 (right). The graph and the Morse indices are known to us, while concerning the numbering we know the following. The lowest vertex in Fig. 7 has number 1, the vertices in the middle row have numbers 2, 3, 4, and 5, and the vertices in the upper row have the numbers 6, 7, and 8. The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5. Indeed, the intersection index $\langle\Delta_{5},\Delta_{6}\rangle$ of cycles vanishing in two newborn Morse critical points is equal to 1, see Lemma V.3.2 in [24]. In particular, the positive critical values of the polynomial $f!$ correspond to all vertices in the top row of Fig. 7 and one vertex of valence three in the middle row.
All virtual functions with intersection matrix defined by the graph of Fig. 7 and the Morse indices and numbering of vertices satisfying these restrictions belong to the same virtual component. Indeed, they can all be obtained from each other by transpositions of orders of vertices that are not connected by edges of the graph and correspond to the critical values of the same sign. Such a transposition is a flip of virtual functions which corresponds to the elementary function surgery when one critical value overtakes another critical value at a distant critical point. This surgery is not related with crossing the discriminant. Our program, when applied to one of the virtual functions from this set, says that this virtual component has 1216 vertices. $\Box$
@@
The set of all polynomials $F(l)$, $l\in L$, is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l). Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin. These numbers $c(l)$ form a lower semi-continuous function on $L/{\mathbb{Z}}_{2}$. By the previous lemma, this function is positive everywhere. Therefore, it is separated from zero by a constant $\delta>0$. The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant. This union is connected and invariant under the action (44) of the group $S(3)$. By Lemma 49, for $l\equiv\frac{x-y}{\sqrt{2}}$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$. The polynomial $F(l)-\delta$ also belongs to this component because all the polynomials F(l)tF(l)-t, where $t\in[\delta,1]$, are non-discriminant. Therefore, the entire component is $S(3)$-invariant. This concludes the proof of Theorem 44 for the virtual component with $\mbox{Card}=1216$.
Components with $\mbox{Card}=1318$ and $844$. Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant variety in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other real critical points with critical value 0. By the versality of the deformation (42), this stratum is a smooth submanifold of codimension four in $\Xi_{2}$. All polynomials $F(l)$ of the form (61) belong to this stratum.
Consider a tubular neighborhood of this stratum in $\Xi_{2}$ fibered into its orthogonal slices diffeomorphic to small four-dimensional balls centered at the points of the stratum. Suppose this tubular neighborhood is narrow enough that, for any $l\in L$, the fiber over the point $\{F(l)\}$ of the stratum is the parameter space of a versal deformation of the singularity of the function $F(l)$ at the origin. Then, all the intersections of these fibers with the discriminant are locally ambient isotopic to the discriminant set of the model versal deformation
(62) $X^{3}-3XY^{2}+Z^{2}+\varkappa_{1}+\varkappa_{2}X+\varkappa_{3}Y+\varkappa_{4}(X^{2}+Y^{2})$
of the sample polynomial $X^{3}-3XY^{2}+Z^{2}$ of class $D_{4}^{-}$ with parameters $\varkappa_{i}$.
Definition 12.
For each $l\in L$, denote the fiber of this tubular neighborhood over the point $\{F(l)\}$ by $U(l)$, and define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
@@
    $f$ has four real critical points near the origin that are obtained by splitting the critical point of the polynomial $F(l)$: specifically, one minimum and three critical points with Morse index 1,
@@
Similarly, define the subset $W(l)\subset U(l)$ as the union of polynomials $f$ satisfying condition (a) but with condition (b) replaced by $\mbox{Card}(f)=1318$.
Remark 51.
According to Proposition 43 and Lemma 50, the set $V(l)\cup W(l)$ is the union of all polynomials in $U(l)$ with $\mbox{Ind}=-1$ that satisfy condition (a).
@@
Preliminary remarks. By the versality property, the discriminant is equisingular along the smooth stratum $\{D_{4}^{-}\}$. Therefore, it is sufficient to prove Lemma 52 for only one particular $l\in L$, say, for $l\equiv\frac{x-y}{\sqrt{2}}$.
@@
To prove Lemma 52, we will describe the geometry of the discriminant in the parameter space of the versal deformation (62) in detail. The Morse polynomials of the form (62) can have two or four critical points depending on the parameters $\varkappa_{2},\varkappa_{3},$ and $\varkappa_{4}$. In a neighborhood of the origin of the space ${\mathbb{R}}^{3}$ of these parameters, the set of non-Morse polynomials looks as shown in Fig. 11 (left), see [3], § 21.3.
@@
This set divides the space ${\mathbb{R}}^{3}$ into three parts. The larger part consists of polynomials with only two real critical points. Two symmetric “pyramids” are filled with polynomials (62) that have three critical points with Morse index 1 and one critical point with Morse index 0 or 2, depending on the pyramid. We are only interested in the polynomials (62) with a minimum point. For each generic interior point $(\varkappa_{2},\varkappa_{3},\varkappa_{4})$ of the corresponding pyramid, the set of non-discriminant polynomials (62) with these values or parameters consists of five intervals. Indeed, the coefficient $\varkappa_{1}$ of a polynomial from this set should be not equal to minus a critical value of the polynomial
(63) $X^{3}-3XY^{2}+Z^{2}+\varkappa_{2}X+\varkappa_{3}Y+\varkappa_{4}(X^{2}+Y^{2}),$
@@
i.e., it can belong to one of the five intervals into which these critical values divide the line $\mathbb{R}$}. For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal. The set of these non-generic polynomials (63) divides the pyramid into six parts. The intersection of the pyramid with a plane orthogonal to its axis looks as the curvilinear triangle shown in Fig. 11 (center). The non-generic polynomials form three segments in it. In particular, the central point of this triangle is represented by the polynomial of the form (63) with $\varkappa_{2}=\varkappa_{3}=0$ and equal critical values at all three points with Morse index 1. These critical values are equal to $\frac{4}{27}\varkappa_{4}^{3}$. The orthogonal plane sections of the pyramid are defined by the conditions $\varkappa_{4}=C$ for different constants $C>0$. The set of all polynomials (62) having a minimum point and three critical points with critical value zero consists of one curve in ${\mathbb{R}}^{4}$ satisfying the conditions
(64) $\varkappa_{2}=\varkappa_{3}=0,\quad\varkappa_{1}=-\frac{4}{27}\varkappa_{4}^{3},\quad\varkappa_{4}>0.$
The zero-level sets of all such polynomials in the plane $\{Z=0\}$ look as in Fig. 10 (right).
@@
Choose an arbitrary polynomial inside the pyramid and label its three critical points of index 1 as $a,b,$ and $c$. This labeling of critical points can be uniquely extended by continuity to all other polynomials within the pyramid. Each of the six domains, into which the curvilinear triangle in Fig. 11 (center) is split by the segments, is characterized by the order of the critical values of the corresponding polynomials at the critical points labeled in this way.